2. The work done by substance in a reversible non-flow manner in accordance with V=700/P where V is

volume in m 3 , P is absolute pressure in Kpaa. Evaluate the work done on or by the substance as the

pressure increases from 85 to 810 KPa


1
Expert's answer
2021-09-15T11:37:53-0400

"W = \int PdV \\

V= \frac{700}{P} \\

W = (-\frac{700}{P^2})dP \\

W = \int P(\frac{-700}{P^2}dP"

"W = -700 \int \frac{dP}{P}" from 85 to 810 kPa

"W = -700(ln(810) -ln(85)) \\

W = -700ln(\frac{810}{85}) = -700ln(9.529) \\

W = -700 \times 2.254 = -1577.8 \;kN-m"